Instructions: Use the data tool to answer the questions below. Choose one, two or several series as directed and press Plot Reset in between activities. Economic Data (1960- 2017) ECONOMIC DATA (1960-2017) Recessions Off Legend 24% 20% 16% 12% 8% 4% 0% -4% 1980:0 1960 1965 1970 1975 1980 1985 1990 1995 2000 2005 2010 2015 1980 1960 2017 ※ SETTINGS □Moodys Baa LT Rate Real GDP Growth Rate of Inflation □ Unemployment Rate □ M1 Money Growth 1M2 Money Growth Fed Funds Nominal Fed Funds Real Fed Funds Target 3 Month T-bill Rate 1 Year T-note Rate 2 Year T-note Rate 5 Year T-note Rate 10 Year T-note Rate 30 Year T-bond Rate Moodys Aaa LT Rate Risk Spread Tenn spread Ex-post Real Rate Instructions: Select the real GDP growth rate and unemployment rate series. Turn on Recessions at the top of the graph to add recession bars to the graph

Saved 0% -4% 1960 1965 1970 1975 1980 1985 1990 1995 2000 2005 2010 2015 1980 1960 2017 SETTINGS Olot Ecaph Reset Real GDP Growth Rate of Inflation Unemployment Rate M1 Money Growth M2 Money Growth Fed Funds Nominal Fed Funds Real Fed Funds Target 3 Month T-bil Rate 1 Year T-note Rate 2 Year T-note Rate 5 Year T-note Rate 10 Year T-note Rate 30 Year T-bond Rate □Moodys Aaa LT Rate Moodys Baa LT Rate Risk Spread Term Spread EEx-post Real Rate Instructions: Select the real GDP growth rate and unemployment rate series. Turn on Recessions at the top of the graph to add recession bars to the graph a) Look at the graph of the GDP growth rate and unemployment series. Focusing on the 2007-2009 recession, check all that apply from the following list of statements. The unemployment rate falls immediately at the end of a recession. The GDP growth rate falls during a period of recession. The unemployment rate rises during a recession The GDP growth may become negative during a recession.

Suppose that furniture production encompasses the following stages: Stage 1 Trees are sold to lumber company Stage 2: Lumber is sold to furniture company. Stage 3: Furniture company sells furniture to retail store Stage 4 Furniture store sells furniture to consumer $1,000 $2500 $6,000 $12,000 Instructions: Enter your responses rounded to the nearest whole number a. What is the value added at each stage? Stage 1 $ Stage 2 S Stage 3: $ Stage 4 S b. How much does this output contribute to GDP? C. How much would the output contribute to GDP if the lumber were imported from Canada?

b. gross domestic product. gross national product. national income. c. d. Use the data below to answer questions 3-6: Consumption Expenditures (C) $850 billion Government Expenditures (G) $300 billion Investment Expenditures ()$275 billion Exports (EX) Imports (IM) Depreciation $50 billion $75 billion S30 billion Gross Domestic Product (GDP) is equal to a. $1,550 billion b. $1,450 billion c. $1,400 billion d. $1,375 billion Net Domestic Product (NDP) is equal to a. $1,420 billion b. $1,350 billion 4. S1,525 billion $1,370 billion d. 5. Net investment expenditures equal a. $245 billion b. $275 billion c. $305 billion d. $30 billion Net exports are equal to a. $50 billion b. $25 billion C. d. $25 billion $75 billion
Answer #1

Value added= Value of production of firm - value of intermediate goods used by the firm

[A] Stage 1 - 1000

Stage 2 - 1500 [2500-1000]

Stage 3 - 3500 [6000-2500]

Stage 4 - 6000 [12000- 6000]

[B] GDP = Sum total of value added

so contribution to GDP = 12000

[C] If the lumber were imported from Canada then contribution to GDP = 12000 - 1000 = 11000, because this 1000 is national income of Canada.   

[3-6]

3. GDP = C+I+G+[EX- IM]

850 + 275 + 300 + [50 - 75]

1425 - 25 = 1400

4. NDP = GDP - Depreciation

1400 - 30 = 1370

5. Net Investment = Gross Investment - depreciation

= 275 - 30 = 245

6. Net exports = exports - imports

= 50 - 75 = -25